OpenStudy (anonymous):

Who can help me with this question: f ' (x) = (x+3)(2x-5). Find f (x) when f ' (x) is given.

sam (.sam.):

Expand and integrate

OpenStudy (anonymous):

can anoyone help me and guide me to do this problem bcoz what my teacher explained I didnt get it.

sam (.sam.):

Pretty straight forward for this one, \[f'(x)=2 x^2+x-15\] \[f(x)=\int\limits 2x^2+x-15 dx\]

sam (.sam.):

\[f(x)=\frac{2}{3}x^3+\frac{x^2}{2}-15x+c\]

sam (.sam.):

Use power rule for each terms, \[\int\limits x^n~dx= \frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

isnt that \[\frac{ x ^{2} }2\] is \[\frac{ x ^{3} }{ 2 }\]

sam (.sam.):

no, you only add one for the exponent

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so to use the power rule, so am I going to plug in 2x?

sam (.sam.):